pure silver can be obtained by heating Ag2CO3. given 25g of Ag2CO3. 
a. how many grams of ag is produced?
b. how many molecules of co3 are produced?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Unlike other metal carbonates that usually decomposes into metal oxides liberating carbon dioxide, silver carbonate on heating decomposes into elemental silver liberating mixture of carbon dioxide and oxygen gas as Ag2CO3 (s) 2Ag (s) + CO2 (g) + 1/2 O2 (g) MW = 276g 2 × 108 = 216 g Hence, 25 g of Ag2CO3 on heating will give 216/276× 25 = 19.56 g Ag as residue.